Calculate the [OH-], pH, and pOH if the [H3O+] = 2.7x10-7 M
1
Expert's answer
2020-05-06T13:58:49-0400

Solution:

The pH is calculated using the expression:

pH = - log[H3O+]

pH = - log(2.7×10-7) = 6.568 = 6.57

pH = 6.57


The pH and pOH of a water solution at 25oC are related by the following equation:

pH + pOH = 14.

Then,

pOH = 14 - pH

pOH = 14 - 6.57 = 7.43

pOH = 7.43


The pOH is calculated using the expression:

pOH = - log[OH-]

[OH-] = 10-pOH

[OH-] = 10-7.43 = 3.7×10-8

[OH-] = 3.7×10-8 M


Answer:

pH = 6.57

pOH = 7.43

[OH-] = 3.7×10-8 M